OpenStudy (anonymous):

The temperature of a pan of hot water varies according to Newton's Law of Cooling: dT/ dt equals negative k times the quantity T minus A, where T is the water temperature, A is the room temperature, and k is a positive constant. If the water cools from 90°C to 85°C in 1 minute at a room temperature of 30°C, how long (to the nearest minute) will it take the water to cool to 60°C? a)3 b)4 c)5 d)8

OpenStudy (anonymous):

@ganeshie8 @Luigi0210 @Destinymasha

OpenStudy (anonymous):

@rational can u plz help? :)

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@Ashleyisakitty can u plz help? :)

OpenStudy (anonymous):

@nirmalnema

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[90-30=60\] and \[85-30=55\] so you can model this as \[T(t)=60\times \left(\frac{55}{60}\right)^t\] where\(T(t)\) is the difference in temperatures

OpenStudy (anonymous):

or even \[T(t)=60\times \left(\frac{11}{12}\right)^t\]

OpenStudy (anonymous):

when the water is 60 degrees it will be \(60-30=30\) degrees hotter than the room solve \[30=60\times \left(\frac{11}{12}\right)^t\] using logs

OpenStudy (anonymous):

so it would be: 1/2 = (11/12)^t ?:)

OpenStudy (anonymous):

the solution is: t=log(2)/log(11/12)

OpenStudy (anonymous):

i got it, its 8 :) thank u!!!!!